### v1

# Text-Tuple-Table: Towards Information Integration in Text-to-Table Generation via Global Tuple Extraction

Abstract
The task of condensing large chunks of textual information into concise and structured tables has gained attention recently due to the emergence of Large Language Models (LLMs) and their potential benefit for downstream tasks, such as text summarization and text mining. Previous approaches often generate tables that directly replicate information from the text, limiting their applicability in broader contexts, as text-to-table generation in real-life scenarios necessitates information extraction, reasoning, and integration. However, there is a lack of both datasets and methodologies towards this task. In this paper, we introduce LiveSum, a new benchmark dataset created for generating summary tables of competitions based on real-time commentary texts. We evaluate the performances of state-of-the-art LLMs on this task in both fine-tuning and zero-shot settings, and additionally propose a novel pipeline called T3(Text-Tuple-Table) to improve their performances. Extensive experimental results demonstrate that LLMs still struggle with this task even after fine-tuning, while our approach can offer substantial performance gains without explicit training. Further analyses demonstrate that our method exhibits strong generalization abilities, surpassing previous approaches on several other text-to-table datasets. Our code and data can be found at https://github.com/HKUST-KnowComp/LiveSum-TTT.
[…]
tcb@breakable
Text-Tuple-Table: Towards Information Integration in Text-to-Table Generation via Global Tuple Extraction
Zheye Deng, Chunkit Chan, Weiqi Wang, Yuxi $\varheartsuit\varheartsuit{}^{\varheartsuit}$, Wei Fan, Tianshi Zheng, Yauwai Yim, Yangqiu Song Department of Computer Science and Engineering, HKUST, Hong Kong SAR, China $\varheartsuit\varheartsuit{}^{\varheartsuit}$chool of Computer Science, Fudan University, Shanghai, China {zdengah, yqsong}@cse.ust.hk
1 Introduction
Figure 1: An overview of the differences between our proposed LiveSum dataset and previous dataset Wiseman et al. (2017), as well as our proposed pipeline called T3(Text-Tuple-Table) which consists of three steps.
Reading extensive texts is demanding and time-consuming for humans, further compounded by the challenge of effectively capturing the key elements. Consequently, recent works have shifted to explore the structured summarization of text Jain et al. (2024), with tables being one highly prevalent form Wu et al. (2022); Li et al. (2023c); Sundar et al. (2024). These approaches improve text comprehension by extracting inherent yet valuable structural information from long unstructured text and enabling their applications in downstream scenarios, such as question answering Chen et al. (2020); Zhu et al. (2024), text summarization Wiseman et al. (2017); Wang et al. (2020); Mulwad et al. (2023) and text data mining Li et al. (2023b); Sui et al. (2024).
However, previous studies on text-to-table generation primarily rely on datasets traditionally used for table-to-text tasks Wiseman et al. (2017); Novikova et al. (2017). One evident issue is that these tasks focus merely on format transformation, where the information in the table and the corresponding text representation are essentially similar Lebret et al. (2016); Bao et al. (2018). For example, in the upper part of Figure 1, the table can be easily completed by extracting relevant numbers from the text without intermediate inference. Such seemingly meaningful correlations can introduce bias into the models, causing them to excel at replicating relevant information but struggle when it comes to categorizing and integrating numbers in complex scenarios. This is further evidenced by the fact that fine-tuning models already perform very well and greatly surpass zero-shot LLMs Tang et al. (2023); Sundar et al. (2024). Hence, a more complicated dataset that requires information aggregation and minimizes the presence of spurious correlations, closely resembling real-world scenarios, is definitely needed for a more rigorous evaluation of models’ text-to-table generation proficiency.
Apart from the research gap at the benchmark level, in terms of methodology, considerable attention has been given to studying the ability of LLMs to comprehend and generate complex structured outputs Tang et al. (2023); Jain et al. (2024), driven by the exceptional success of LLMs in recent years Touvron et al. (2023); Anthropic (2024); OpenAI (2024). Extensive benchmarks indicate that LLMs exhibit sub-optimal performance in zero-shot settings, with multiple cases of generating inaccurate contents deviate from the given text Tang et al. (2023). To address this issue, more sophisticated prompting mechanisms have been proposed Wei et al. (2022); Khot et al. (2022); Dua et al. (2022). Among them, Jain et al. (2024) introduce a divide-and-generate prompting approach to generate more accurate and informative tables, demonstrating its effectiveness in improving model performance. However, this simplistic approach of dividing text into paragraphs and generating tables is unsuitable in more complicated situations because table-relevant information may not be contiguous in the original text and may span across various paragraphs. Therefore, developing a robust prompting method is also needed for generating complex tables that capture crucial information from scattered text or paragraphs.
To resolve the aforementioned research gaps, we introduce a novel benchmark, LiveSum, which consists of 3,771 text-based live commentaries from real-world football matches, intending to evaluate the models’ ability to generate summary tables. Unlike previous benchmarks, our benchmark necessitates the model to possess the ability to extract correct and meaningful information from complex textual data, specifically emphasizing information integration, reasoning, and conceptualization skills Wang et al. (2023a). This is because commentaries in close temporal proximity or with similar semantic meanings may describe the same event, while verbs with similar meanings may refer to the same types of events. For example, in Figure 1, the second and third dialogue boxes both describe the same goal event, and the verbs “goal” and “score” refer to the same goal event.
Along with the benchmark, we also introduce a robust prompting-based method T3 to address our proposed task. Specifically, our method draws inspiration from the inherent attributes of the table, where each cell, along with its corresponding row header and column header, creates an informative triple (namely (row header, column header, cell)), which degenerates into a binary tuple when lacking row or column headers. These tuples serve as cues for humans to locate specific information in the text and complete the table accordingly. Consequently, our pipeline begins by extracting the relevant tuples from the text, followed by the integration of these tuples, and ultimately generating one or more summary tables.
We hope that the proposed dataset, method, and experimental results can provide valuable insights for tasks such as text-to-table, as well as any task involving the generation of complex structured outputs from text. In summary, in this paper, we make the following contributions:
  • To the best of our knowledge, LiveSum is the first benchmark dataset designed to evaluate the information integration ability of models in text-to-table generation tasks.
  • We introduce a novel T3(Text-Tuple-Table) prompting pipeline that functions as a flexible framework, applicable to any text-to-table generation tasks.
  • We conduct extensive experiments to evaluate the performance of LLMs under different settings and demonstrate that our T3 pipeline can bring significant improvements while showcasing excellent generalization capabilities.
2 Task Definition
We first provide a formal definition of the text-to-table generation. The input $\mathcal{S}$ consists of a textual passage with $n$ tokens, denoted as $\mathbf{x}=x_{1},\dots,x_{n}$, and optionally, an instruction text with $m$ tokens, denoted as $\mathbf{y}=y_{1},\dots,y_{m}$, which provides guidance on the format or content of the generated tables. The output $\mathcal{T}$ is a set of $k(k\geq 1)$ tables, $\mathbf{T}^{1},\dots,\mathbf{T}^{k}$. For the output tables, we present a more detailed definition that covers two aspects: structure-related and content-related.
Structure
We assume there are no merged cells in the tables for simplicity. Each table $\mathbf{T}^{i}$ has a caption $\mathbf{c}^{i}=c^{i}_{1},\dots,c^{i}_{d}$, where $d=|\mathbf{c}^{i}|$, and it consists of $n^{i}_{r}$ rows and $n^{i}_{c}$ columns, resulting in a total of $n^{i}_{r}\times n^{i}_{c}$ cells. The cell $\mathbf{T}^{i}_{p,q}$ in the $p$-th row and $q$-th column is composed of a sequence of tokens: $\mathbf{T}^{i}_{p,q,1},\dots,\mathbf{T}^{i}_{p,q,r}$, where $r=|\mathbf{T}^{i}_{p,q}|$. The table $\mathbf{T}^{i}$ must have either a row header for all rows or a column header for all columns and it is also possible for the table to have both.
Content
We define that the information in the output tables should be derived from the input $\mathbf{x}$ or can be inferred from $\mathbf{x}$. For each cell $\mathbf{T}^{i}_{p,q}$ in table $\mathbf{T}^{i}$, when combined with its row header $\mathbf{T}^{i}_{p,1}$ (if any), its column header $\mathbf{T}^{i}_{1,q}$ (if any), and the table’s caption $\mathbf{c}^{i}$, it should convey the equivalent information as expressed in the input $\mathbf{x}$ and complies with the instruction $\mathbf{y}$ (if any).
3 LiveSum Dataset
Figure 2: Overview of the pipeline for constructing the LiveSum dataset illustrated with a sample sentence.
We consider the problem of generating match statistic information tables from textual live commentary. Inspired by RotoWire Wiseman et al. (2017), a data-to-document dataset in the sports domain that aims to generate textual summaries by incorporating statistical data from basketball games, we instead focus on live commentary in football, which is available on BBC Sports [^1] https://www.bbc.com/sport/football. We crawl the data for the English Premier League from 2014 to 2023 and obtain complete commentary for 3,771 matches. Figure 2 shows the pipeline we use to construct the dataset. Section 3.1 describes the generation process of the live commentary, and Section 3.2 describes the generation process for the summary table. More details are provided in Appendix A.
3.1 Live Commentary Generation
To address formatting issues in the original textual live commentary on the website, we paraphrase the text to match the commentator’s style while ensuring a certain degree of diversity. Building upon previous studies Kim et al. (2023a); Chen et al. (2023); Kim et al. (2023b), we employ ChatGPT OpenAI (2022) to generate complete live commentary automatically. Subsequently, to comply with privacy regulations and prevent bias in LLM benchmarks, we also anonymize the data leveraging named entity recognition (NER) techniques Qi et al. (2020) to produce the final version of the live commentary.
Figure 3: Eight types of event information (inner circle) that require summarization in LiveSum dataset, along with their common expressions (outer circle) in the commentary.
3.2 Summary Table Generation
On the other hand, human annotators label a summary table for each match’s commentary. We recruit five workers who are interested in football and are from English-speaking countries to perform the annotations. Since the occurrence of the events in football matches is deterministic, the ground truth is essentially unambiguous. In cases where there are inconsistencies in the annotated results, the correct answer is determined through a majority vote.
3.3 Statistics
LiveSum comprises a collection of 3,771 pairs, consisting of textual live commentaries and corresponding summary tables. We randomly split the entire dataset into training and test sets, resulting in 3,017 instances for the training set and 754 instances for the test set. On average, each live commentary segment consists of 1,256 words. LiveSum focuses on eight types of events, with the names and their corresponding common descriptions displayed in Figure 3.
4 T3(Text-Tuple-Table) Pipeline
Our proposed T3(Text-Tuple-Table) pipeline is designed to mimic the intuitive steps followed by humans when performing this task. When individuals aim to summarize a table from text, they typically extract pertinent or valuable tuples from the content, guided by any provided instructions, and then organize these tuples into one or more tables. Based on this concept, we divide this transformation into three stages: text-to-tuple, integration, and tuple-to-table, each of which is discussed in the following subsections. Taking Figure 1 as an example, we first extract key events mentioned in the text, then aggregate this information into consolidated tuples, and ultimately compile them into a table.
4.1 Text-to-Tuple
[…]
Figure 4: The performance of various LLMs under fine-tune and zero-shot settings, as well as after the application of the T3 method on the test set of LiveSum dataset. The average RMSE and error rate for each model are displayed, along with the error rate for each of the three difficulty sections. More results are in Table 1.
Considering the superior performance and flexibility of LLMs in information extraction compared to traditional techniques Ma et al. (2023); Xu et al. (2023), we employ an LLM as our tuple extractor. We follow the instructions from InstructUIE Wang et al. (2023b) and design the following prompting:
Text-to-Tuple Prompting Template According to <Instruction>, please extract the relevant events and information in the form of tuples, structured as (subject, object, verb) or (subject, attribute, value): <Text>
where <Instrction> is the directive for the current task, and <Text> is the text to be transformed.
4.2 Information Integration
In this stage, we propose two approaches for integrating information. The first one involves direct execution by the LLM, using prompting to consolidate tuple data. The second one uses algorithms and code generated by the LLM to integrate tuple information, inspired by the LLMs’ great success in code generation tasks Roziere et al. (2023); Luo et al. (2023); Guo et al. (2024). T3 defaults to using code generation in this step. The promptings for these two methods are shown as follows:
Information Integration Prompting Template Direct Execution: According to <Instruction>, please integrate these tuples as required: <Tuples> Code Generation: According to <Instruction>, please develop an algorithm to consolidate these tuples as specified: <Tuples>
where <Instrction> is the directive for the current task, and <Tuples> consists of the tuples extracted in the prior stage.
4.3 Tuple-to-Table
After obtaining the integrated tuples, we follow the previous implementation Tang et al. (2023); Jain et al. (2024) and use the following prompting to generate the final tables:
Tuple-to-Table Prompting Template According to <Instruction>, please generate one or more tables based on the following tuples: <Tuples>
where <Instrction> is the directive for the current task, and <Tuples> consists of the tuples produced in the prior stage.
5 Experimental Setup
Baseline Models
In this study, we conduct fine-tuning on the LiveSum dataset using three representative open-source LLMs: Mistral-7B-Instruct-v0.2 Jiang et al. (2023), LLaMA-2 Chat 7B and LLaMA-2 Chat 13B Touvron et al. (2023). We fine-tune these models following the current state-of-the-art fine-tuning methodologies Tang et al. (2023). Therefore, the outcomes represent the best results achievable with the present fine-tuning methods. We also evaluate eight state-of-the-art LLMs in zero-shot settings: LLaMA-2 Chat 13B, LLaMA-2 Chat 70B Touvron et al. (2023), Mistral Large MistralAI (2024), Claude 2.1 Anthropic (2023), Claude 3 Opus Anthropic (2024), ChatGPT OpenAI (2022), and GPT-4 OpenAI (2024). For each model, we conduct tests using two types of prompts. The first type directly describes the task by providing an instruction text $\mathbf{y}$ and accompanying it with the text $\mathbf{x}$. The second type uses the Chain-of-Thought (CoT) prompting Wei et al. (2022), incorporating the phrase “let’s think step by step” into the instruction text. See more details in Appendix C.
Evaluation Metric
As the generated cell content in this task consists of numerical values, we utilize commonly employed metrics in regression tasks, namely the Root Mean Square Error (RMSE). We also report the Error Rate (ER) for each cell, defining a cell as erroneous if its content does not exactly match the ground truth.
Grouping by Event Difficulty
Furthermore, we categorize the eight types of events into three groups based on assessed difficulty: Goals, due to direct descriptions of scores in the original text, and Red Cards, due to their rare occurrence are categorized into the Easy section. Shots and Fouls, due to their varied expressions and descriptions, are classified into the Hard section. The remaining four event types are classified as Medium section. We report the RMSE and ER for each model across different difficulty categories to provide a more comprehensive analysis.
6 Experiments and Analyses
In this section, we will benchmark the performance of current state-of-the-art LLMs on the LiveSum dataset, and further evaluate the effectiveness and generalization of our proposed approach. We aim to answer the following research questions:
RQ1 (Benchmarking)
How do the current state-of-the-art LLMs perform on this dataset in fine-tuning and zero-shot settings?
RQ2 (Effectiveness)
How does our proposed T3 pipeline impact model performance?
RQ3 (Generalization)
How effective is the T3 pipeline when applied to other real-world datasets for the text-to-table generation task?
Model Easy Medium Hard Average
RMSE ER RMSE ER RMSE ER RMSE ER
[…]
Mistral-7B-Instruct-v0.2 1.045 38.36 3.832 85.11 7.115 95.52 4.564 76.03
LLaMA-2-Chat 7B 1.047 38.41 3.728 84.91 7.107 95.40 4.512 75.91
LLaMA-2-Chat 13B 1.043 39.22 3.587 84.60 6.671 94.42 4.287 75.71
[…]
LLaMA-2-Chat 13B 0.775 33.29 4.554 87.37 5.203 93.29 4.279 75.33
LLaMA-2-Chat 13B (CoT) 0.780 31.83 4.376 87.42 5.088 92.35 4.162 74.75
LLaMA-2-Chat 70B 0.410 12.34 3.189 88.59 4.941 92.41 3.455 70.48
LLaMA-2-Chat 70B (CoT) 0.450 12.86 3.221 89.25 5.314 94.24 3.613 71.40
ChatGPT 0.200 8.06 2.864 72.73 4.257 90.62 3.008 61.03
ChatGPT (CoT) 0.229 10.61 2.809 72.75 4.087 90.38 2.911 61.62
Claude 2.1 1.014 10.08 2.581 63.99 4.621 90.58 3.171 57.16
Claude 2.1 (CoT) 1.496 14.06 2.291 61.70 4.081 90.38 2.918 56.96
Mistral Large 0.005 0.27 2.385 52.45 2.712 84.62 2.209 47.45
Mistral Large (CoT) 0.018 0.73 2.311 51.82 2.608 84.08 2.139 47.12
GPT-4 0.156 4.64 1.167 46.05 4.114 88.53 2.273 46.32
GPT-4 (CoT) 0.154 4.38 1.173 45.86 3.981 88.73 2.225 46.20
Claude 3 Opus 0.078 2.52 1.617 51.36 3.713 88.06 2.253 48.33
Claude 3 Opus (CoT) 0.040 1.59 1.642 49.60 3.265 87.86 2.079 47.17
[…]
Claude 2.1 (T3) 0.193 8.95 1.965 44.99 2.751 72.15 2.066 42.77
Mistral Large (T3) 0.191 8.82 1.596 42.37 2.136 69.23 1.631 40.70
GPT-4 (T3) 0.056 3.18 0.854 25.83 1.219 46.22 0.929 25.27
Claude 3 Opus (T3) 0.081 5.30 0.406 14.79 0.477 21.29 0.438 14.04
Table 1: The performance of various LLMs under three settings, showing RMSE and error rate across three difficulty categories and overall average. We bold the best results and underline the second-best results in each setting.
6.1 Benchmarking (RQ1)
We first analyze the performance of existing state-of-the-art LLMs on the LiveSum dataset under fine-tuning and zero-shot settings, with results displayed in Figure 4 and Table 1. Overall, the performance of most models in the zero-shot setting far exceeds that in the fine-tuning setting, indicating that the previous state-of-the-art fine-tuning method has limited capability for information integration, and there is substantial room for improvement on this benchmark. In the zero-shot setting, it is noteworthy that most models show a slight improvement in both metrics after applying CoT. Among them, the best-performing models are Mistral Large, GPT-4, and Claude 3 Opus, which are nearly comparable. They achieve RMSEs ranging from 2.08 to 2.27 and error rates between 46.20% and 48.33%. Nevertheless, this still highlights a notable deficiency in the information integration capabilities of LLMs in the zero-shot settings, underscoring the challenges and significance of our benchmark. We then analyze performance across three categories of difficulty.
Easy Section
It can be observed that the error rate of the fine-tuned models is generally around 40%, with an RMSE close to 1. In the zero-shot setting, LLaMA-2-Chat, ChatGPT, and Claude 2.1 models exhibit relatively poor performance, occasionally producing anomalously large values. The error rates of the other models generally remain below 5%, with RMSEs less than 0.2. Among these, the Mistral Large model performs the best, with both metrics significantly lower than other models.
Medium Section
The medium section exhibits the greatest variation among models and serves as a crucial determinant of overall model performance. We organize the models based on their performance, with error rates in the zero-shot setting ranging from 89.25% down to 45.86%. GPT-4 with CoT performs the best, achieving the lowest error rate, which still indicates the suboptimal capabilities of LLMs.
Hard Section
The charts clearly show that in the hard section, the zero-shot method shows minimal enhancement compared to fine-tuning, as most error rates are around 90%. However, Mistral Large is an exception, achieving a lower error rate of 84.08%, which demonstrates the challenging nature of the hard section.
6.2 Effectiveness (RQ2)
We apply the T3 pipeline to four LLMs: Claude 2.1, Mistral Large, GPT-4, and Claude 3 Opus; the rest are not applicable to this method due to their ineffective extraction of tuples from inputs of such lengths, resulting in a minimal number of tuples or a substantial duplication of the same tuple. The implementation details are discussed in Appendix B. From Figure 4 and Table 1, it is observable that Claude 2.1 and Mistral Large exhibit similar improvements after applying the T3 method, both showing slight enhancements over the best results in the zero-shot setting, with reductions in RMSE of 34.9% and 26.2% respectively and decreases in error rate of 25.2% and 14.2%. In contrast, GPT-4 and Claude 3 Opus display substantial improvements after implementing the T3 method, with RMSE reductions of 59.1% and 80.6%, respectively, and error rate reductions of 45.4% and 70.9%. The reductions in these metrics are primarily reflected in the significant decrease in the error rate for the hard sections, creating a clear distinction from the zero-shot approaches. We then conduct an ablation study to evaluate the impact of the T3 method on model performance. We experiment with two variant methods: T3-Merged (T3m) and T3-Direct-Execution (T3d). The former method involves using a single prompt that instructs the model to first extract relevant tuples before generating the table, while the latter one modifies the second step of the T3 method to be directly executed by the LLM, rather than using code generation (see Appendix B.5 for more details). Table 2 presents the results using GPT-4 as an example. It is apparent that using T3 considerably enhances the model’s overall performance, leading significantly in overall metrics, with average reductions in RMSE and error rate of 59.1% and 45.4%, respectively. Compared to GPT-4 and CoT, the variants T3m and T3d also exhibit notable improvements, with respective reductions in RMSE of 28.7% and 37.6%, and error rates of 8.6% and 10.3%. Full ablation studies are in Appendix D.
6.3 Generalization (RQ3)
Model Easy Medium Hard Average
RMSE/ER RMSE/ER RMSE/ER RMSE/ER
GPT-4 0.16/4.6 1.17/46.1 4.11/88.5 2.27/46.3
w/ CoT 0.15/4.4 1.17/45.9 3.98/88.7 2.23/46.2
w/ T3m 0.00/0.1 1.42/43.2 2.46/82.8 1.62/42.3
w/ T3d 0.09/4.5 1.12/40.1 2.23/81.4 1.42/41.5
w/ T3 0.06/3.2 0.85/25.8 1.22/46.2 0.93/25.3
Table 2: The ablation study results comparing the performance of different prompting methods. We bold the best results and underline the second-best results.
To examine the generalization capabilities of T3, we apply T3 to two additional datasets designed to test text-to-table performance and compare it with previous methods. Section 6.3.1 involves table generation without the need for information integration, while Section 6.3.2 focuses on table generation without instructions.
6.3.1 Performance of T3 on Struc-Bench Table Dataset
Metric Zero-Shot Fine-Tune
ChatGPT w/ T3 LLaMA-7B
SacreBLEU 77.58 78.91 90.60
ROUGE-L 86.11 88.36 88.98
BERTScore 96.75 97.34 98.54
BLEURT 64.66 67.47 66.07
BARTScore  -2.08  -1.90  -0.69
Content P-Score   6.84    7.29   7.69
Format P-Score   9.70    9.88   8.60
Content H-Score   1.66    1.68   1.65
Format H-Score   3.28   3.63   3.61
Table 3: The evaluation results on the test set of Struc-Bench Table dataset with nine metrics. We bold the best results and underline the second-best results.
We test the performance of T3 on the text-to-table benchmark Struc-Bench Table Tang et al. (2023). This benchmark is based on the Rotowire dataset Wiseman et al. (2017) and employs traditional evaluation metrics, prompting score(P-Score), and heuristical score(H-Score), to conduct a comprehensive assessment of the output tables. They also introduce a fine-tuning approach incorporating row and column header information in the training instructions. We argue that this comparison with the zero-shot method is problematic. In zero-shot settings, due to the absence of a fine-tuning process, the format of the output tables is uncertain. For example, in ground-truth tables, cells that are left blank may be filled with terms such as “unknown” or “not mentioned” by the model, substantially impacting similarity-based metrics. Hence we modify the model’s outputs under zero-shot settings before reporting the results. We evaluate the performance of ChatGPT with and without employing the T3 method and also compare it to the fine-tuned LLaMA-7B model proposed by Tang et al. (2023). Results are detailed in Table 3. It is observable that the application of the T3 method results in significant improvements across all metrics, with some measures outperforming the fine-tuned model. Further details on the experiments and analysis of the results are discussed in Appendix E.
6.3.2 Performance of T3 on Wiki40b Dataset
We intend to evaluate the performance of our proposed approach in the text-to-table task without instructions and ground-truth tables. In line with the pioneering work of StructSum Jain et al. (2024), we experiment on a randomly sampled set from the English section of Wiki40b dataset Guo et al. (2020). As there is no ground-truth table for the text in the dataset, they leverage LLMs and propose Auto-QA Coverage as an evaluation metric:
$\text{Cov}(\mathcal{T})=\frac{\sum_{i=1}^{|G(\mathcal{S})|}E_{(q_{i},a_{i})}[Q(\mathcal{T},q_{i})]}{|G(\mathcal{S})|}$
where $G(\mathcal{S})$ is the list of Question-Answer pairs $(q_{i},a_{i})$ generated by the LLM based on text $\mathcal{S}$, $Q(\mathcal{T},q)$ is the LLM’s answer to question $q$ based on table $\mathcal{T}$, and $E_{(q,a)}[x]$ is the LLM’s evaluation of whether answer $a$ and $x$ are equivalent for question $q$. On top of this evaluation, we add a step where an LLM is used to pre-screen each $(q,a)$ pair based on text $\mathcal{S}$, filtering out any pairs where the question can not be correctly answered. This process further assures the quality of the QA pairs generated by $G$. We opt for ChatGPT as the LLM for evaluation and randomly sample 500 passages for the test dataset following Jain et al. (2024). We also introduce T2(Text-Tuple) which treats the intermediary tuples from T3 as a single table $\mathcal{T}$. We aim to investigate the extent of information loss during the conversion from tuples to tables through this configuration. Figure 5 shows the Auto-QA Coverage of three methods. The curve indicates the percentage of generated tables meeting a given coverage threshold. Overall, T3 demonstrates a substantial improvement over the prior Devide-and-Generate method Jain et al. (2024). For example, when the coverage threshold is set to 70%, about 65% of data reach this threshold after applying T3, compared to only 50% with the preceding approach. It is important to note that T2 outperforms T3, with 83% of data retaining the same coverage after tuple extraction. This suggests that information loss occurs during the transformation of raw tuples into structured tables. Exploring ways to mitigate this loss represents an essential area for further research. More details are discussed in Appendix F.
00101010102020202030303030404040405050505060606060707070708080808090909090100100100100101010102020202030303030404040405050505060606060707070708080808090909090100100100100Percentage (%)Auto-QA Coverage (%)Divide-and-Generate Jain et al. (2024) T3(Text-Tuple-Table) T2(Text-Tuple)
Figure 5: Auto-QA coverage of the three methods. The point $(P,C)$ means $P\%$ of the data can achieve a coverage of $C\%$ or higher measured using Auto-QA.
6.4 Case Studies
We present specific case studies on the outputs of different models on the LiveSum dataset in Appendix G. These cases directly demonstrate the effectiveness of our proposed method. Additionally, we summarize some common errors of the model after applying T3 and areas for improvement.
7 Related Work
Text-to-Table Generation
Many studies have been proposed to perform text-to-table generation, converting it into sequence-to-sequence problems Wu et al. (2022); Li et al. (2023c), or framing them as question-answering problems Sundar et al. (2024). With the rise of LLMs, some research has also explored evaluating LLMs under fine-tuning or zero-shot settings, and it shows that fine-tuning yields highly effective results Tang et al. (2023); Sundar et al. (2024). However, these methods employ datasets that only require the model to extract relevant information from text and populate tables, which significantly limits the scope of this task. Therefore, we introduce a new challenging dataset and propose a universal solution that greatly enhances the performance of LLMs under zero-shot setting.
LLMs for Information Extraction
Information Extraction (IE) is critical and foundational for many downstream tasks in NLP. Many works have been conducted to leverage LLMs and provide effective solutions for IE tasks within a generative framework Ma et al. (2023); Lu et al. (2023); Wan et al. (2023); Zhou et al. (2024). Recent progress in LLMs also has led to the development of unified frameworks that model various IE tasks and domains Wang et al. (2023b); Sainz et al. (2024). This aligns with our intention to harness this capability to address general text-to-table generation tasks.
LLM Promptings
Prompt engineering has been essential for enhancing LLMs and has demonstrated great success across a wide range of applications Wei et al. (2022); Dua et al. (2022); Li et al. (2023a); Wang et al. (2024). Among the various prompting techniques, we find the decomposed prompting Khot et al. (2022), which breaks down complex tasks into easier sub-tasks via prompting, highly effective for text-to-table generation. Jain et al. (2024) adopts this idea by breaking the text into small pieces for table generation. However, we argue that such a decomposition approach is impractical because text is not always easily divisible. For example, in our dataset, such division might result in adjacent sections describing the same event, causing errors. Hence we propose a more intuitive and broadly applicable task decomposition pipeline.
8 Conclusion
In this work, we introduce LiveSum, a novel and challenging benchmark dataset for assessing the capability of models to integrate information in the text-to-table generation, along with a robust pipeline named T3. Experimental results show that current LLMs underperform on our dataset in both fine-tuning and zero-shot settings; however, significant improvements are observed after applying our proposed T3 pipeline. Our method can also be applied to any text-to-table dataset, enabling LLMs to outperform previous methods in zero-shot settings.
Limitations
Although our LiveSum benchmark extensively evaluates the information integration capabilities of LLMs, we have not yet tested their performance in a few-shot setting. Despite the challenge posed by the token length of live commentary for few-shot settings, we reserve this aspect for future work. Furthermore, while our proposed T3 pipeline significantly improves performance on several state-of-the-art LLMs, it cannot be effectively applied to LLMs that are deficient in tuple extraction capabilities, as it fails in the first stage and cannot proceed to the next phase. Developing methods that boost performance on such LLMs remains a valuable area for future research.
Ethics Statement
When constructing the LiveSum dataset, we sample texts of live football match commentary from the open-access BBC Sports official website. We apply LLMs to paraphrase this live commentary and conduct manual reviews to ensure no harmful content is generated. We also anonymize the data using named entity recognition (NER) technology combined with player rosters. The datasets used in our experiments, Struc-Bench Tang et al. (2023) and Wiki40b Guo et al. (2020), are open-source, and all experiments adhere to their intended use for research purposes. Therefore, to the best of the authors’ knowledge, we believe that this work introduces no additional risk.
Acknowledgements
The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20), and the GRF (16211520 and 16205322) from RGC of Hong Kong. We also thank the UGC Research Matching Grants (RMGS20EG01-D, RMGS20CR11, RMGS20CR12, RMGS20EG19, RMGS20EG21, RMGS23CR05, RMGS23EG08).
Appendices
Appendix A Details of LiveSum Dataset
This section extends Section 3, where we discuss additional details of dataset construction, and present detailed statistical information and a specific example of the dataset.
A.1 Live Commentary Generation
Paraphrasing
We first utilize ChatGPT to paraphrase the original text to make it closer to the language used by the commentator while ensuring a certain degree of diversity. The prompting template used in this step is as follows:
Paraphrase Prompting Template You are a football commentator, paraphrase the sentence and don’t make it too long: <Original Text>
where <Original Text> represents the original live commentary on the webpage. After the transcription of each live commentary segment, we manually inspect whether the meaning of each sentence has been altered. If any issues are identified, we make manual revisions accordingly.
Data Anonymization
We then perform the data anonymization. First, we obtain the list of players participating in each match from the BBC Sports official website. Then, using a combination of string matching and NER techniques Qi et al. (2020), we match fully detected names to players on the list based on textual similarity. Each individual is assigned a unique number and recorded in the format Player<number>(<team>), where <number> is a positive integer and <team> is the team that player belongs to. We anonymize the team names as “Home Team” and “Away Team”. The inclusion of team names is intended to eliminate the need for the model to infer the team affiliation of each player, thus avoiding any potential interference with our evaluation of their information integration ability. We also manually inspect whether there is any missed anonymization. If any are found, we make manual revisions accordingly.
A.2 Summary Table Generation
Annotation
We recruit five workers who are interested in football and are from English-speaking countries to annotate the events and employ a majority voting approach to resolve any disagreements. We randomly sample 100 cases with discrepancies and find that all errors are due to carelessness. Therefore, we conclude that the quality of the annotation quality is assured.
Integration
After manual annotation and review, the data are aggregated into several tuples, and we count the occurrence of each tuple to obtain the integrated result. Finally, we sequentially fill in the column headers of the table with the eight events, convert the team names to Home Team and Away Team for the row headers of the table, and then fill in the corresponding counts into the cells to generate the final summary table.
A.3 Statistics
Table 4 presents the statistical data regarding text length and the frequency of various events within the train and test sets of the LiveSum dataset. It can be seen that the statistical metrics across the train set and test set are comparably uniform, consistent with our method of completely random division.
Train Set Test Set
# Instances 3,017    754
Average Text Length
Words 1,258 1,250
Chars 6,852 6,816
Average Event Occurrence Frequency
Goals   1.38   1.39
Shots 12.71 12.55
Fouls 10.60 10.57
Yellow Cards   1.74   1.76
Red Cards   0.04   0.03
Corner Kicks   5.25   5.24
Free Kicks 10.34 10.30
Offsides   1.86   1.84
Table 4: Average statistics in LiveSum dataset.
A.4 Example
Figure 6 presents an example of the live commentary and a summary table generated through the pipeline depicted in Figure 2.
Textual Live Commentary Players are being announced for the lineup and getting ready for the game. The game is now underway with the start of the first half. Offside called on Home Team as Player5(Home Team) attempts a through ball to Player11(Home Team), who is caught offside. Player28(Away Team) makes a goal with a right-footed shot from the center of the box, giving the Away Team a 1-0 lead over the Home Team. Player8(Home Team) earns a free kick on the right side of the field. Player23(Away Team) commits a foul. Player9(Home Team) commits a foul. Player27(Away Team) earns a free kick in their own half. The Home Team earns a corner kick. Player5(Home Team)’s left footed shot from outside the box is saved in the bottom left corner after an assist from Player7(Home Team). Player26(Away Team) scores with a right-footed shot from outside the box, assisted by Player29(Away Team), Home Team 0, Away Team 2. The Home Team wins a corner kick. Player8(Home Team) commits a foul. Player21(Away Team) earns a free kick in their own half. Player29(Away Team) scores with a right-footed shot from the right side of the box, assisted by Player28(Away Team), Home Team 0, Away Team 3. Player7(Home Team) of the Home Team attempts a through ball, but Player11(Home Team) is flagged for being offside. Player29(Away Team) misses the goal with a high right footed shot from outside the box, assisted by Player26(Away Team). The Home Team wins a corner kick. Player10(Home Team) attempts a through ball, but Player11(Home Team) is offside for the Home Team. The Away Team wins a corner kick. Player5(Home Team) earns a free kick in their own half. Player21(Away Team) commits a foul. Player25(Away Team) fouls Player5(Home Team), who earns a free kick on the left wing. The Home Team wins a corner kick. Player10(Home Team) of the Home Team is caught offside after Player8(Home Team) attempts a through ball. At the end of the first half, the Home Team is trailing with a score of 0-3 against the Away Team. And we’re back for the second half with the Home Team trailing 0-3 against the Away Team. Player2(Home Team) earns a free kick in their own half. Player23(Away Team) commits a foul. Player18(Home Team)’s header from the center of the box is saved in the bottom left corner. Player4(Home Team) commits a foul. Player29(Away Team) earns a free kick in their own half. Player4(Home Team) has received a yellow card for a reckless foul. Player29(Away Team) is currently delayed in the match due to an injury. The delay is finished and they are prepared to resume play. Player8(Home Team) is holding up the game due to an injury. The delay is finished and they are prepared to resume play. Player14(Home Team) commits a foul. Player23(Away Team) earns a free kick in the opponent’s half. Player28(Away Team) missed the target with a shot from the right side of the box, with an assist from Player27(Away Team) after a quick counterattack. Player28(Away Team)’s right footed shot from the centre of the box was close, but missed to the right and then blocked. Player29(Away Team)’s shot from the center of the box is saved in the center of the goal after an assist from Player28(Away Team). The Away Team earns a corner kick. Player28(Away Team)’s shot from the center of the box is blocked with the assistance of Player23(Away Team). Player26(Away Team) attempts a through ball, but Player23(Away Team) is offside for the Away Team. Player14(Home Team) from the Home Team attempts a through ball, but Player11(Home Team) is flagged for being offside. The Home Team wins a corner kick. Player6(Home Team)’s close-range attempt is saved in the bottom left corner with an assist from Player3(Home Team)’s headed pass. Player18(Home Team) earns a free kick on the left wing after being fouled by Player20(Away Team). Player26(Away Team) attempts a through ball but Player20(Away Team) is offside for the Away Team. Player4(Home Team) commits a foul. Player28(Away Team) earns a free kick in their own half. Player27(Away Team)’s shot from the center of the box was blocked with the help of Player26(Away Team). Player6(Home Team) is being delayed in the match due to an injury. The delay is finished and they are prepared to resume play. Player18(Home Team)’s shot from outside the box with their right foot was too high, assisted by Player5(Home Team). Player6(Home Team) commits a foul on Player23(Away Team), resulting in a free kick being awarded on the left wing. The Away Team gets a corner kick. Player23(Away Team)’s left footed shot from outside the box following a corner is too high. Player25(Away Team) attempted a shot with his right foot from the right side of the box but it was too high, with an assist from Player28(Away Team). Player7(Home Team) earns a free kick in the opponent’s half. Player36(Away Team) commits a foul. Player3(Home Team) misses the header to the left from the center of the box, assisted by Player16(Home Team) with a cross. Player28(Away Team) misses the target with a left-footed shot from long range. Player16(Home Team) attempts a through ball, but Player18(Home Team) is offside for the Home Team. The Home Team finishes the second half with a score of 0, while the Away Team has 3. Final score, Home Team 0, Away Team 3. Summary Table Team Goals Shots Fouls Yellow Cards Red Cards Corner Kicks Free Kicks Offsides Away Team 3 12 6 0 0 3 6 2 Home Team 0 5 6 1 0 5 6 6
Figure 6: An instance within the LiveSum dataset. The above section presents a complete textual live commentary that has undergone paraphrasing and data anonymization, while the table below represents a manually annotated summary table.
Appendix B Implementation of T3 on LiveSum Dataset
This section is an extension of Section 4, where we provide description of the implementation details of the T3 method applied to the LiveSum dataset.
B.1 Instrction on LiveSum Dataset
We present the instruction directives common to all experiments conducted on the LiveSum dataset, which are some rules related to this task:
Instruction on LiveSum Dataset According to the live text, please count the number of: 1.goals, 2. shots, 3.fouls, 4.yellow cards, 5.red cards, 6.corner kicks, 7.free kicks, and 8.offsides for each team. Note that goals and saved attempts and blocked attempts and missed attempts are considered shots. Handball and dangerous play are also considered foul. The second yellow card is also considered a red card. Penalty is also considered as free kicks.
B.2 Text-to-Tuple Prompts
Based on the characteristics of the text in the Livesum dataset, we have defined the format of tuples as (player name, team name, event) or (team name, event). Subsequently, we define the following prompting template:
T3
Text-to-Tuple Prompting Template
<Instruction> Please extract all the relevant event from the following passage, output them in (player name, team name, event) or (team name, event) format. Constrain the event names to only the following options: 1.goals, 2.shots, 3.fouls, 4.yellow cards, 5.red cards, 6.corner kicks, 7.free kicks, and 8.offsides: <Text>
where <Instruction> represents the text provided in Appendix B.1, and <text> is the live commentary text.
B.3 Information Integration Prompts
As discussed in Section 4.2, T3 employs code generation as the default approach for information integration. The following is the prompting template we define, and Figure 7 presents an example of Python code generated by GPT-4.
T3
Information Integration Prompting Template
Code Generation: <Instruction> Please develop a Python code to consolidate these tuples as specified: <Tuples>
where <Instruction> represents the text provided in Appendix B.1, and <Tuples> is the tuples extracted in the previous step.
# Initialize a dictionary to store the counts
team_stats = {}
# List of tuples
tuples = [(’player1’, ’hometeam’, ’goal’),
(’player2’, ’hometeam’, ’shot’),
(’player3’, ’awayteam’, ’foul’)]
# Loop through the tuples
for player, team, event in tuples:
# If the team is not in the dictionary, add it
if team not in team_stats:
team_stats[team] = {’goals’: 0, ’shots’: 0, ’fouls’: 0, ’yellowcards’: 0, ’redcards’: 0,
’cornerkicks’: 0, ’freekicks’: 0, ’offsides’: 0}
# Update the counts based on the event
if event == ’goal’:
team_stats[team][’goals’] += 1
team_stats[team][’shots’] += 1
elif event in [’shot’, ’savedattempt’, ’blockedattempt’, ’missedattempt’]:
team_stats[team][’shots’] += 1
elif event in [’foul’, ’handball’, ’dangerousplay’]:
team_stats[team][’fouls’] += 1
elif event == ’yellowcard’:
team_stats[team][’yellowcards’] += 1
elif event in [’redcard’, ’secondyellowcard’]:
team_stats[team][’redcards’] += 1
if event == ’secondyellowcard’:
team_stats[team][’yellowcards’] += 1
elif event == ’cornerkick’:
team_stats[team][’cornerkicks’] += 1
elif event in [’freekick’, ’penalty’]:
team_stats[team][’freekicks’] += 1
elif event == ’offside’:
team_stats[team][’offsides’] += 1
# Print the counts for each team
for team, stats in team_stats.items():
print(f"{team}:{stats}")
Figure 7: Python code generated by GPT-4 in response to the prompting as described in Appendix B.3.
B.4 Tuple-to-Table Prompts
In this step, we present the format of the generated table and specify the following prompting template:
T3
Tuple-to-Table Prompting Template
<Instruction> Please only output a table with the team name in CSV format with 2 rows based on the following tuples: <Tuples>
where <Instruction> represents the text provided in Appendix B.1, and <Tuples> is the tuples integrated by the code in the previous step.
B.5 Prompts of Variant Methods
In Section 6.2, two variants of the T3 method are mentioned during the design of the ablation study. Here, we present the implementation details of these variants.
T3-Merged
This method combines all three stages of T3 into one, resulting in a single prompt:
T3-Merged
Prompting Template
<Instruction> Let’s do the following things: 1. Extract all the relevant events from the following passage in (player name, team name, event) or (team name, event) format. 2. Integrate these tuples. 3. Output a table with 2 rows in CSV format. <Text>
where <Instruction> represents the text provided in Appendix B.1, and <text> is the live commentary text.
T3-Direct-Execution
This method replaces code generation with direct execution in the second stage of T3. The prompts are as follows:
T3-Direct
Information Integration Prompting Template
<Instruction> Please count all the information required and integrate these tuples: <Tuples>
where <Instruction> represents the text provided in Appendix B.1, and <Tuples> is the tuples extracted in the previous step. The first and third steps remain unchanged and are described exactly as in Appendix B.2 and  B.4.
Appendix C Details of Benchmark Configuration
This section serves as an extension of Section 5, providing additional details regarding the benchmark configuration.
C.1 Fine-Tune Setting
We use the open-source library named LLaMA-Factory [^2] https://github.com/hiyouga/LLaMA-Factory Zheng et al. (2024) to fine-tune all models. LoRA Hu et al. (2022) is used as the fine-tuning paradigm to accommodate our computational resources. The pre-trained weights are downloaded from the huggingface library Wolf et al. (2020). We load the models with FP16 as the precision and optimize them with an Adam optimizer Kingma and Ba (2015). The learning rate is set to 5e-5 and the batch size is 32. The maximum length for the input and generated sentence concatenation is 3,500. We warm up the model with 3,000 steps and evaluate the model every 300 steps. A linear scheduler is also used. The LoRA rank is set to 8, and the $\alpha$ is set to 32.
C.2 Zero-Shot Setting
The zero-shot and chain-of-thought inference for LLaMA-2 models are conducted on eight V100 GPUs. When obtaining outputs from LLMs via APIs, we ensure deterministic results by setting the temperature to zero. In Table 5, we list the model names corresponding to different model types. Note that unless specifically stated otherwise, when a model type from the table is used, the model name corresponding to it in the table is the one being invoked.
Model Type Model Name
Mistral Large mistral-large-2402
ChatGPT gpt-3.5-turbo-0125
GPT-4 gpt-4-0613
Claude 2.1 claude-2.1
Claude 3 Opus claude-3-opus-20240229
Table 5: Model types and their corresponding names of LLMs used in our experiments.
C.3 Prompts
Below are the two prompting templates we used to evaluate the baseline models:
Baseline Prompting Template Prompt w/o CoT <Instruction> Please output a table with 2 rows in CSV format according to the following live text: <Text>
Baseline Prompting Template Prompt w/ CoT <Instruction> Let’s think step by step! At last, please output a table with 2 rows in CSV format according to the following live text: <Text>
where <Instruction> represents the text provided in Appendix B.1, and <text> is the live commentary text.
C.4 Result Parsing
As the table format in this experiment is fixed, most models can produce recognizable tables. We employ a comprehensive evaluation algorithm based on regular expressions to parse various types of output. For improperly formatted tables in the model’s output, we filtered out those instances. This is because when the table format is disrupted, we are unable to obtain the meaning of each number in the output, and thus cannot calculate the metric. Moreover, this indicates that the model is even unable to perform the text-to-table generation task, which falls outside the scope of assessing its information integration abilities in the text-to-table process.
Appendix D Full Results of Ablation Study
Model Easy Medium Hard Average
RMSE ER RMSE ER RMSE ER RMSE ER $\Delta_{\text{ER}}$
Claude 2.1 1.014 10.08 2.581 63.99 4.621 90.58 3.171 57.16
Claude 2.1 (CoT) 1.496 14.06 2.291 61.70 4.081 90.38 2.918 56.96 $\downarrow$   0.3%
Claude 2.1 (T3m) 1.144 13.92 2.289 62.28 4.114 90.12 2.869 57.15 $\downarrow$   0.0%
Claude 2.1 (T3d) 3.653 39.15 2.444 64.88 5.621 92.06 4.056 65.24 $\uparrow$ 14.1%
Claude 2.1 (T3) 0.193 8.95 1.965 44.99 2.751 72.14 2.066 42.77 $\downarrow$ 25.2%
Mistral Large 0.005 0.27 2.385 52.45 2.712 84.62 2.209 47.44
Mistral Large (CoT) 0.018 0.73 2.311 51.82 2.608 84.08 2.139 47.12 $\downarrow$   0.7%
Mistral Large (T3m) 0.039 1.84 2.223 52.70 3.479 86.18 2.399 48.36 $\uparrow$   1.9%
Mistral Large (T3d) 0.142 6.59 1.677 57.46 2.735 84.81 1.865 51.58 $\uparrow$   8.7%
Mistral Large (T3) 0.191 8.82 1.596 42.37 2.137 69.23 1.631 40.70 $\downarrow$ 14.2%
GPT-4 0.156 4.64 1.167 46.05 4.114 88.53 2.273 46.32
GPT-4 (CoT) 0.154 4.38 1.163 45.86 3.981 88.73 2.225 46.20 $\downarrow$   0.3%
GPT-4 (T3m) 0.003 0.13 1.419 43.22 2.458 82.76 1.621 42.34 $\downarrow$   8.6%
GPT-4 (T3d) 0.087 4.47 1.124 40.13 232 81.45 1.418 41.55 $\downarrow$ 10.3%
GPT-4 (T3) 0.056 3.18 0.854 25.83 1.219 46.22 0.929 25.27 $\downarrow$ 45.4%
Claude 3 Opus 0.078 2.52 1.617 51.36 3.713 88.06 2.253 48.33
Claude 3 Opus (CoT) 0.040 1.59 1.642 49.60 3.265 87.87 2.079 47.17 $\downarrow$   2.4%
Claude 3 Opus (T3m) 1.244 15.14 1.610 52.97 3.625 85.27 2.426 51.59 $\uparrow$   6.7%
Claude 3 Opus (T3d) 0.327 9.66 1.315 46.43 1.924 79.50 1.432 45.50 $\downarrow$   5.9%
Claude 3 Opus (T3) 0.081 5.30 0.406 14.79 0.477 21.29 0.438 14.04 $\downarrow$ 70.9%
Table 6: The comprehensive ablation study results comparing the performance of different prompting methods. We bold the best results and underline the second-best results.
In this section, we extend Section 6.2 and provide a comprehensive supplementary ablation study. We apply two variant methods, T3-Merged (T3m) and T3-Direct-Execution (T3d), proposed in Section 6.2, to all four LLMs capable of applying the T3 method. The results of all experiments are listed in Table 6. In this table, we also provide a separate listing for each model, showcasing the average change in error rate for each method compared to the baseline. It is important to note that a lower error rate indicates better performance.
Impact of CoT Prompting
Firstly, we observe that CoT prompting yields positive effects across all four LLMs, resulting in an overall reduction in error rate ranging from 0.3% to 2.4%. This further corroborates that CoT is a concise and effective prompting strategy.
Impact of T3m Prompting
After applying the T3m method, it is observed that the performance of the Mistral Large and Claude 3 Opus deteriorates, particularly with a significant increase in error rate of 6.7% for Claude 3 Opus. This indicates that the models’ ability to summarize the information internally is superior to the approach of extracting and merging information separately. On the contrary, GPT-4 demonstrates a significant reduction in error rate of 8.6% after applying T3m, which indicates that the capabilities of GPT-4 are sufficiently robust to support the T3m prompting approach and effectively execute all three steps. The change in error rate for Claude 2.1 is not significant; however, there is a notable 9.6% reduction in RMSE. This indicates that T3m leads to a closer approximation of the true values in its results.
Impact of T3d Prompting
After applying T3d, the error rates of Claude 2.1 and Mistral Large increase by 14.1% and 8.7%, respectively, indicating that they are not suitable for this prompting method. Comparing these results to those obtained with T3, it can be inferred that the reason for the increase in error rate is likely due to their inferior ability to integrate information compared to the generated code, as the error rate significantly decreases when using code integration. On the other hand, GPT-4 and Claude 3 Opus achieved error rate reductions of 10.3% and 5.9%, respectively, under this prompting approach. This indicates that both models possess a certain level of ability to integrate information.
Impact of T3 Prompting
All four LLMs showed substantial improvements after applying the T3 prompting method. Specifically, Mistral Large, Claude 2.1, and GPT-4 achieve average error rate reductions of 14.2%, 25.2%, and 45.4% respectively. Notably, Claude 3 Opus exhibits a remarkable error rate reduction of 70.9%. This strongly indicates the effectiveness of the proposed method. From the perspective of absolute error rates, under the T3 method, the performance of the four models from best to worst is as follows: Claude 3 Opus, GPT-4, Mistral Large, and Claude 2.1. This ranking essentially reflects their capabilities in tuple extraction.
Appendix E Details of the Experiment on Struct-Bench Table Dataset
This section is an extension of Section 6.3.1, in which we will introduce the evaluation criteria for this dataset, our implementation details, and an analysis of the results.
Model SacreBLEU ROUGE-L BERTScore BLEURT BARTScore Content P-Score Format P-Score Content H-Score Format H-Score
[…]
ChatGPT 77.58 86.11 96.75 64.66 -2.08 6.84 9.70 1.66 3.28
ChatGPT (T3) 78.91 88.36 97.34 67.47 -1.90 7.29 9.88 1.68 3.63
GPT-4 87.26 92.81 98.15 77.08 -1.58 7.45 9.71 1.76 3.87
GPT-4 (T3) 87.75 92.37 98.30 79.80 -1.60 7.60 9.74 1.77 3.89
[…]
LLaMA-7B 90.60 88.98 98.54 66.07 -0.69 7.69 8.60 1.65 3.61
Table 7: Evaluation results on the test set of Struc-Bench Table dataset, consisting of nine metrics. We bold the best results and underline the second-best results.
E.1 Evaluation Metrics
This benchmark employs nine evaluation metrics, five of which are traditional and applicable to text generation tasks: SacreBLEU Post (2018), ROUGE-L Lin (2004), BERTScore Zhang et al. (2020), BLEURT Sellam et al. (2020), and BARTScore Yuan et al. (2021). These metrics can to some extent measure the similarity between the generated tables and the target tables. The last four metrics, P-score (Prompting Score) and H-score (Heuristical Score), are evaluation criteria proposed by Tang et al. (2023), which involve using ChatGPT to score the content and format of generated tables and applying manually devised rules to assess the content and format, respectively. For all evaluation metrics, higher numbers signify better performance.
E.2 Implementation Details
We conduct experiments on Struct-Bench Table dataset Tang et al. (2023) using ChatGPT and GPT-4 to assess the impact of T3 on model performance. As described in Section 6.3.1, we first obtain the outputs of ChatGPT and GPT-4 using the provided prompting, removed terms such as “unknown” and “not mentioned” from the outputs, and then recalculate all metrics. This step involves using all instructions, prompting templates, and the code for calculating all metrics, all of which are sourced from the codebase [^3] https://github.com/gersteinlab/Struc-Bench provided by Tang et al. (2023). We then apply our proposed T3 method to both ChatGPT and GPT-4. We design the prompting template for the first step as follows:
T3 on Struct-Bench Table Dataset
Text-to-Tuple Prompting Template
<Instruction> You are now required to extract team and player information from the following input. Please focus on the table format and extract all relevant tuples in (team or player name, attribute, value) format: <Text>
where <Instruction> is provided from the dataset, and <text> is the original text. Because the content of this dataset does not involve information integration, the second step does not alter the tuples from the first step. Therefore, we proceed directly to the third step of table generation, setting the following prompting template:
T3 on Struct-Bench Table Dataset
Tuple-to-Table Prompting Template
<Instruction> Based on the instructions and the following extracted tuples, please generate two tables according to the table format: <Tuples>
where <Instruction> is provided in the dataset, and <Tuples> is the tuples extracted in the previous step.
E.3 Results Analysis
Table 7 presents all our experimental results, with the fine-tuning section quoting the state-of-the-art results Tang et al. (2023). The results indicate that GPT-4 outperforms ChatGPT across all metrics. After applying the T3 method, both ChatGPT and GPT-4 show improved performance, with ChatGPT experiencing a greater enhancement, suggesting the generalizability of the T3 method. When compared with state-of-the-art fine-tuning methods, the zero-shot approaches perform better on some metrics. This indicates that the T3 method still offers significant improvements for text-to-table tasks that do not require information integration, further demonstrating its broad applicability.
Appendix F Details of the Experiment on Wiki40b Dataset
This section is an extension of Section 6.3.2, in which we will discuss more details.
F.1 Evaluation Metrics
As introduced in Section 6.3.2, the evaluation criterion adopted for this dataset, Auto-QA Coverage, is designed due to the absence of ground truth for the generated tables. Jain et al. (2024) propose using question answering as a medium and leverage LLMs to assess the quality of the generated tables. This metric not only measures how much information from the original text is covered by the table but also checks the accuracy of the values within the table since incorrect content would also lead to errors in the QA. They also demonstrate through detailed experiments that Auto-QA Coverage aligns with human evaluation standards, and thus we follow this metric in our experiments.
F.2 Implementation Details
Since Jain et al. (2024) samples 100 entries from the English version of Wiki40b Guo et al. (2020) without disclosing which ones, we replicate their setting by sampling 500 texts to serve as the dataset for this experiment. For each passage, we utilize ChatGPT to generate 20 (question, answer) pairs that can be answered based on the text. Subsequently, we introduced a verification step where all (question, answer) pairs are re-fed into ChatGPT to confirm their correctness. Pairs that ChatGPT cannot accurately answer are filtered out. This approach enhances the accuracy of metric evaluation. Here, we present the prompts for T3:
T3 on Wiki40b Dataset
Text-to-Tuple Prompting Template
You are going to summarize a table for this passage, but the first step is to extract useful information. Output them in (subject, attribute, value) or (subject, verb, object) format: <Text>
where <text> is the original passage. It is noteworthy that, under this task setting, there are no instructions provided. Next, we feed the output tuples into the second step, where the LLM autonomously integrates the tuples:
T3 on Wiki40b Dataset
Information Integration Prompting Template
Please integrate these tuples if necessary: <Tuples>
Finally, the tuples output from this step are fed into the prompt for the last step.
T3 on Wiki40b Dataset
Tuple-to-Table Prompting Template
Summarize the triples below in one or multiple tables. Use the following format: Caption: A caption for the table you generate. It can be multiple lines. Table: A table in markdown format. <Tuples>
The other prompts used in the experiments, such as those for generating question-answer pairs, assessing the correctness of answers, and the Divide-and-Generate method’s prompting, all originate from Jain et al. (2024).
F.3 Results Analysis
The results of this experiment are presented in Figure 5. As discussed in Section 6.3.2, T3 demonstrates higher Auto-QA Coverage than the baseline, proving its ability to generate higher quality tables in text-to-table generation tasks without instructions. We also experiment with using the tuples extracted in the first step of T3 directly as the generated tables and assessed their coverage, resulting in constructive findings. The coverage curve of T2 is entirely above that of T3, indicating a certain loss of information from tuple to table. Although this could also be due to insufficient table question answering capabilities leading to a decrease in metrics, it is necessary to employ relevant techniques Wu et al. (2023); Wang et al. (2024) to mitigate the impact of this factor in future work. Nevertheless, the conclusion that T3 improves the performance of the previous work remains sound.
Appendix G Case Studies
Figure 8 lists the outputs of four LLMs with and without the application of the T3 method on the data shown in Figure 6. For the results not utilizing T3, we can not easily analyze why it generates a range of large or small values. However, for the results using the T3 method, we perform a detailed examination. We randomly sample 100 results generated by GPT-4 applying T3 and conduct a spot check, finding that all errors originated from the first stage. Among these errors, 78% are due to missing event tuples, and 21% are due to wrong event tuples. Here, we present two representative examples.
Missing Event Tuple Example Player18(Home Team) earns a free kick on the left wing after being fouled by Player20(Away Team). (Player18, Home Team, Free Kick) (Player20, Away Team, Foul)(missing)
Wrong Event Tuple Example Player17(Home Team) from the Home Team draws a foul in the penalty area, resulting in a penalty conceded by Player33(Away Team). (Player17, Home Team, Foul)(wrong) (Player33, Away Team, Foul) (Player17, Home Team, Free Kick)(missing)
Outputs from Claude 2.1 Results w/o T3: RMSE=1.888, Error Rate=43.75%. Team Goals Shots Fouls Yellow Cards Red Cards Corner Kicks Free Kicks Offsides Away Team 3 ✓ 13 ✗ 10 ✗ 0 ✓ 0 ✓ 3 ✓ 5 ✗ 4 ✗ Home Team 0 ✓ 10 ✗ 7 ✗ 1 ✓ 0 ✓ 5 ✓ 6 ✓ 3 ✗ Results w/ T3: RMSE=0.661, Error Rate=25.00%. Team Goals Shots Fouls Yellow Cards Red Cards Corner Kicks Free Kicks Offsides Away Team 3 ✓ 12 ✓ 4 ✗ 0 ✓ 0 ✓ 3 ✓ 6 ✓ 2 ✓ Home Team 0 ✓ 5 ✓ 7 ✗ 1 ✓ 0 ✓ 6 ✗ 6 ✓ 5 ✗
Outputs from Mistral Large Results w/o T3: RMSE=1.458, Error Rate=50.00%. Team Goals Shots Fouls Yellow Cards Red Cards Corner Kicks Free Kicks Offsides Away Team 3 ✓ 10 ✗ 7 ✗ 0 ✓ 0 ✓ 3 ✓ 10 ✗ 2 ✓ Home Team 0 ✓ 8 ✗ 7 ✗ 1 ✓ 0 ✓ 4 ✗ 7 ✗ 5 ✗ Results w/ T3: RMSE=0.612, Error Rate=18.75%. Team Goals Shots Fouls Yellow Cards Red Cards Corner Kicks Free Kicks Offsides Away Team 3 ✓ 12 ✓ 6 ✓ 0 ✓ 0 ✓ 3 ✓ 4 ✗ 2 ✓ Home Team 0 ✓ 5 ✓ 5 ✗ 1 ✓ 0 ✓ 5 ✓ 5 ✗ 6 ✓
Outputs from GPT-4 Results w/o T3: RMSE=1.785, Error Rate=31.25%. Team Goals Shots Fouls Yellow Cards Red Cards Corner Kicks Free Kicks Offsides Away Team 3 ✓ 10 ✗ 6 ✓ 0 ✓ 0 ✓ 3 ✓ 5 ✗ 2 ✓ Home Team 0 ✓ 11 ✗ 9 ✗ 1 ✓ 0 ✓ 5 ✓ 7 ✗ 6 ✓ Results w/ T3: RMSE=0.433, Error Rate=18.75%. Team Goals Shots Fouls Yellow Cards Red Cards Corner Kicks Free Kicks Offsides Away Team 3 ✓ 12 ✓ 5 ✗ 0 ✓ 0 ✓ 3 ✓ 5 ✗ 2 ✓ Home Team 0 ✓ 5 ✓ 6 ✓ 1 ✓ 0 ✓ 5 ✓ 5 ✗ 6 ✓
Outputs from Claude 3 Opus Results w/o T3: RMSE=2.046, Error Rate=56.25%. Team Goals Shots Fouls Yellow Cards Red Cards Corner Kicks Free Kicks Offsides Away Team 3 ✓ 18 ✗ 7 ✗ 0 ✓ 0 ✓ 2 ✗ 7 ✗ 3 ✗ Home Team 0 ✓ 9 ✗ 7 ✗ 1 ✓ 0 ✓ 5 ✓ 9 ✗ 5 ✗ Results w/ T3: RMSE=0.000, Error Rate=0.00%. Team Goals Shots Fouls Yellow Cards Red Cards Corner Kicks Free Kicks Offsides Away Team 3 ✓ 12 ✓ 6 ✓ 0 ✓ 0 ✓ 3 ✓ 6 ✓ 2 ✓ Home Team 0 ✓ 5 ✓ 6 ✓ 1 ✓ 0 ✓ 5 ✓ 6 ✓ 6 ✓
Figure 8: Case study analysis showing the outputs and evaluation metrics of Claude 2.1, Mistral Large, GPT-4, and Claude 3 Opus with and without the T3 method on data shown in Figure 6.

### v2

@@
 Further analyses demonstrate that our method exhibits strong generalization abilities, surpassing previous approaches on several other text-to-table datasets.
-Our code and data can be found at https://github.com/HKUST-KnowComp/LiveSum-TTT.
+Our code and data can be found at https://github.com/HKUST-KnowComp/LiveSum.
 […]
@@
 (2022); Li et al.
-(2023c); Sundar et al.
+(2023d); Sundar et al.
 (2024).
@@
 (2023) and text data mining Li et al.
-(2023b); Sui et al.
+(2023c); Sui et al.
 (2024).
@@
 (2024), driven by the exceptional success of LLMs in recent years Touvron et al.
-(2023); Anthropic (2024); OpenAI (2024).
+(2023); Anthropic (2024); OpenAI (2024a); Jiang et al.
+(2023b).
 Extensive benchmarks indicate that LLMs exhibit sub-optimal performance in zero-shot settings, with multiple cases of generating inaccurate contents deviate from the given text Tang et al.
@@
 Unlike previous benchmarks, our benchmark necessitates the model to possess the ability to extract correct and meaningful information from complex textual data, specifically emphasizing information integration, reasoning, and conceptualization skills Wang et al.
+(2023b, a, 2024b, 2024c); Wang and Song (2024); Zheng et al.
 (2024a).
@@
 Building upon previous studies Kim et al.
-(2023a); Chen et al.
-(2023); Kim et al.
+(2023); Chen et al.
+(2023); Jiayang et al.
+(2024); Deng et al.
+(2023); Jiayang et al.
+(2023); Chan et al.
 (2024b), we employ ChatGPT OpenAI (2022) to generate complete live commentary automatically.
-Subsequently, to comply with privacy regulations and prevent bias in LLM benchmarks, we also anonymize the data leveraging named entity recognition (NER) techniques Qi et al.
+Subsequently, to comply with privacy regulations and prevent bias in LLM benchmarks Li et al.
+(2023a); Susnjak (2022); Li et al.
+(2024b); Lukas et al.
+(2023); Li et al.
+(2024a), we also anonymize the data leveraging named entity recognition (NER) techniques Qi et al.
 (2020) to produce the final version of the live commentary.
@@
 We follow the instructions from InstructUIE Wang et al.
-(2023b) and design the following prompting:
+(2023c) and design the following prompting:
 Text-to-Tuple Prompting Template According to <Instruction>, please extract the relevant events and information in the form of tuples, structured as (subject, object, verb) or (subject, attribute, value): <Text>
@@
 Information Integration Prompting Template Direct Execution: According to <Instruction>, please integrate these tuples as required: <Tuples> Code Generation: According to <Instruction>, please develop an algorithm to consolidate these tuples as specified: <Tuples>
-where <Instrction> is the directive for the current task, and <Tuples> consists of the tuples extracted in the prior stage.
+where <Instruction> is the directive for the current task, and <Tuples> consists of the tuples extracted in the prior stage.
 4.3 Tuple-to-Table
@@
 Tuple-to-Table Prompting Template According to <Instruction>, please generate one or more tables based on the following tuples: <Tuples>
-where <Instrction> is the directive for the current task, and <Tuples> consists of the tuples produced in the prior stage.
+where <Instruction> is the directive for the current task, and <Tuples> consists of the tuples produced in the prior stage.
 5 Experimental Setup
@@
 In this study, we conduct fine-tuning on the LiveSum dataset using three representative open-source LLMs: Mistral-7B-Instruct-v0.2 Jiang et al.
-(2023), LLaMA-2 Chat 7B and LLaMA-2 Chat 13B Touvron et al.
+(2023a), LLaMA-2 Chat 7B and LLaMA-2 Chat 13B Touvron et al.
 (2023).
@@
 We also evaluate eight state-of-the-art LLMs in zero-shot settings: LLaMA-2 Chat 13B, LLaMA-2 Chat 70B Touvron et al.
-(2023), Mistral Large MistralAI (2024), Claude 2.1 Anthropic (2023), Claude 3 Opus Anthropic (2024), ChatGPT OpenAI (2022), and GPT-4 OpenAI (2024).
+(2023), Mistral Large MistralAI (2024), Claude 2.1 Anthropic (2023), Claude 3 Opus Anthropic (2024), ChatGPT OpenAI (2022), and GPT-4 OpenAI (2024a).
+Because of the limitation of the input lengths, we evaluate two LLMs in few-shot settings: ChatGPT-16k and GPT-4o OpenAI (2024b).
 For each model, we conduct tests using two types of prompts.
@@
 As the generated cell content in this task consists of numerical values, we utilize commonly employed metrics in regression tasks, namely the Root Mean Square Error (RMSE).
+$RMSE=\sqrt{\frac{\sum_{i=1}^{n}\left(y_{i}-\tilde{y}_{i}\right)^{2}}{n}}$
+where $n$ represents the number of cells, $y_{i}$ and $\tilde{y}_{i}$ represent the contents of the cell at index $i$ (for 2D tables, the index is calculated after reshaping to the 1D table) in the ground truth table and the output table, respectively.
 We also report the Error Rate (ER) for each cell, defining a cell as erroneous if its content does not exactly match the ground truth.
@@
 We bold the best results and underline the second-best results in each setting.
+Model Easy Medium Hard Average
+RMSE/ER RMSE/ER RMSE/ER RMSE/ER
+[…]
+0-shot 0.22/12.0 1.76/65.1 3.22/87.2 2.08/57.3
+\ \ $\Diamond$ w/ CoT 0.23/12.7 1.75/64.4 3.14/87.4 2.05/57.2
+1-shot 0.25/13.5 2.29/71.9 4.33/90.4 2.76/61.9
+\ \ $\Diamond$ w/ CoT 0.18/8.82 3.18/73.8 4.52/91.5 3.24/62.0
+[…]
+0-shot 0.00/0.13 1.51/50.4 1.82/78.5 1.43/44.8
+\ \ $\Diamond$ w/ CoT 0.00/0.20 1.31/41.7 2.00/78.1 1.41/40.4
+1-shot 0.00/0.19 1.24/49.0 2.19/78.8 1.44/44.3
+\ \ $\Diamond$ w/ CoT 0.00/0.13 1.19/45.0 2.72/84.3 1.63/43.6
+5-shots 0.00/0.20 1.71/57.2 2.15/80.0 1.65/48.6
+\ \ $\Diamond$ w/ CoT 0.00/0.13 1.55/56.3 2.02/72.9 1.53/46.4
+Table 2: The performance of two LLMs under the few-shot setting. We bold the best results and underline the second-best results.
 6.1 Benchmarking (RQ1)
-We first analyze the performance of existing state-of-the-art LLMs on the LiveSum dataset under fine-tuning and zero-shot settings, with results displayed in Figure 4 and Table 1.
+We first analyze the performance of existing state-of-the-art LLMs on the LiveSum dataset under fine-tuning, zero-shot and few-shot settings, with results displayed in Figure 4, Table 1 and Table 2.
 Overall, the performance of most models in the zero-shot setting far exceeds that in the fine-tuning setting, indicating that the previous state-of-the-art fine-tuning method has limited capability for information integration, and there is substantial room for improvement on this benchmark.
@@
 Nevertheless, this still highlights a notable deficiency in the information integration capabilities of LLMs in the zero-shot settings, underscoring the challenges and significance of our benchmark.
+The results in Table 2 indicate that few-shot learning does not consistently improve model performance on this task.
+However, there is a significant reduction in the error rate for the hard portion under the 5-shots-CoT setting.
+We speculate that in such complex tasks, the model may not effectively learn and understand the examples, and instead, be influenced by the examples, leading to incorrect answers.
 We then analyze performance across three categories of difficulty.
@@
 6.2 Effectiveness (RQ2)
+Model Easy Medium Hard Average
+RMSE/ER RMSE/ER RMSE/ER RMSE/ER
+GPT-4 0.16/4.6 1.17/46.1 4.11/88.5 2.27/46.3
+w/ CoT 0.15/4.4 1.17/45.9 3.98/88.7 2.23/46.2
+w/ T3m 0.00/0.1 1.42/43.2 2.46/82.8 1.62/42.3
+w/ T3d 0.09/4.5 1.12/40.1 2.23/81.4 1.42/41.5
+w/ T3 0.06/3.2 0.85/25.8 1.22/46.2 0.93/25.3
+Table 3: The ablation study results comparing the performance of different prompting methods. We bold the best results and underline the second-best results.
 We apply the T3 pipeline to four LLMs: Claude 2.1, Mistral Large, GPT-4, and Claude 3 Opus; the rest are not applicable to this method due to their ineffective extraction of tuples from inputs of such lengths, resulting in a minimal number of tuples or a substantial duplication of the same tuple.
@@
 6.3 Generalization (RQ3)
-Model Easy Medium Hard Average
-RMSE/ER RMSE/ER RMSE/ER RMSE/ER
-GPT-4 0.16/4.6 1.17/46.1 4.11/88.5 2.27/46.3
-w/ CoT 0.15/4.4 1.17/45.9 3.98/88.7 2.23/46.2
-w/ T3m 0.00/0.1 1.42/43.2 2.46/82.8 1.62/42.3
-w/ T3d 0.09/4.5 1.12/40.1 2.23/81.4 1.42/41.5
-w/ T3 0.06/3.2 0.85/25.8 1.22/46.2 0.93/25.3
-Table 2: The ablation study results comparing the performance of different prompting methods. We bold the best results and underline the second-best results.
 To examine the generalization capabilities of T3, we apply T3 to two additional datasets designed to test text-to-table performance and compare it with previous methods.
@@
 The curve indicates the percentage of generated tables meeting a given coverage threshold.
-Overall, T3 demonstrates a substantial improvement over the prior Devide-and-Generate method Jain et al.
+Overall, T3 demonstrates a substantial improvement over the prior Divide-and-Generate method Jain et al.
 (2024).
@@
 (2022); Li et al.
-(2023c), or framing them as question-answering problems Sundar et al.
+(2023d), or framing them as question-answering problems Sundar et al.
 (2024).
@@
 Recent progress in LLMs also has led to the development of unified frameworks that model various IE tasks and domains Wang et al.
-(2023b); Sainz et al.
+(2023c); Sainz et al.
 (2024).
@@
 (2022); Li et al.
-(2023a); Wang et al.
-(2024).
+(2023b); Chan et al.
+(2023b); Wang et al.
+(2024d); Yim et al.
+(2024); Chan et al.
+(2023a); Lin et al.
+(2024); Chan et al.
+(2024a).
 Among the various prompting techniques, we find the decomposed prompting Khot et al.
@@
 The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20), and the GRF (16211520 and 16205322) from RGC of Hong Kong.
-We also thank the UGC Research Matching Grants (RMGS20EG01-D, RMGS20CR11, RMGS20CR12, RMGS20EG19, RMGS20EG21, RMGS23CR05, RMGS23EG08).
 Appendices
@@
 Annotation
-We recruit five workers who are interested in football and are from English-speaking countries to annotate the events and employ a majority voting approach to resolve any disagreements.
+We recruit five workers who voluntarily agreed to participate in this research without receiving any compensation to annotate the events and employ a majority voting approach to resolve any disagreements.
+These five workers are all postgraduate students who have .
 We randomly sample 100 cases with discrepancies and find that all errors are due to carelessness.
@@
 We use the open-source library named LLaMA-Factory [^2] https://github.com/hiyouga/LLaMA-Factory Zheng et al.
-(2024) to fine-tune all models.
+(2024b) to fine-tune all models.
 LoRA Hu et al.
@@
 C.2 Zero-Shot Setting
-The zero-shot and chain-of-thought inference for LLaMA-2 models are conducted on eight V100 GPUs.
+The zero-shot and chain-of-thought inference for LLaMA-2 models are conducted on eight V100 GPUs Wang et al.
+(2024a).
 When obtaining outputs from LLMs via APIs, we ensure deterministic results by setting the temperature to zero.
@@
 Claude 3 Opus claude-3-opus-20240229
-Table 5: Model types and their corresponding names of LLMs used in our experiments.
-C.3 Prompts
+Table 6: Model types and their corresponding names of LLMs used in our experiments under zero-shot setting.
+C.3 Few-Shot Setting
+Because the input lengths of LiveSum dataset generally exceed 2,000 tokens, even testing with 1-shot would surpass the maximum token limit (4,096) of most LLMs.
+Therefore, we select two models, ChatGPT-16k OpenAI (2022) and GPT-4o OpenAI (2024b), for supplementary experiments Regarding the selection of few-shot examples, we randomly choose samples from the training set where all events occur.
+In the Chain of Thought setting, the reasoning process is included in the examples.
+Model Type Model Name
+ChatGPT-16k gpt-3.5-turbo-16k-0613
+GPT-4o gpt-4o-2024-05-13
+Table 7: Model types and their corresponding names of LLMs used in our experiments under few-shot setting.
+C.4 Prompts
 Below are the two prompting templates we used to evaluate the baseline models:
@@
 where <Instruction> represents the text provided in Appendix B.1, and <text> is the live commentary text.
-C.4 Result Parsing
+C.5 Result Parsing
 As the table format in this experiment is fixed, most models can produce recognizable tables.
@@
 (2023); Wang et al.
-(2024) to mitigate the impact of this factor in future work.
+(2024d) to mitigate the impact of this factor in future work.
 Nevertheless, the conclusion that T3 improves the performance of the previous work remains sound.
+F.4 Error Analysis
+We carry out a supplementary case study on the Wiki40B dataset. We sample 50 articles and observe the following situations:
+  • In the first phase, the average accuracy of the tuples obtained from the text-to-tuple process is 81.6%.
+We define erroneous tuples as those containing: (i) direct extractions of original sentences, (ii) incomplete triples, and (iii) duplicate triples.
+  • In the second phase, tuple integration, 24% of the articles undergo integration of similar items, while the remainder simply involves sorting and categorizing the tuples or outputting them unchanged.
+  • In the third phase, tuple-to-table essentially involves treating each tuple as a row to be filled into a table, but two scenarios may introduce errors: (i) 14% of the articles involve the fragmentation of long sentences, leaving only incomplete tuples, and (ii) 6% of the articles continue to merge similar items based on the tuples, combining related tuples into long sentences.
+(While this step may not necessarily introduce errors, it could increase the difficulty of reasoning, and the AutoQA metric may decrease.)
 Appendix G Case Studies